Problem 10-10 Multiple Products, Materials, and Processes [LO10-1, LO10-2] Mickley Corporation produces two products, Alpha6s and Zeta7s, which pass through two operations, Sintering and Finishing. Each of the products uses two raw materials, X442 and Y661. The company uses a standard cost system, with the following standards for each product (on a per unit basis): Raw Material Standard Labor Time Product X442 Y661 Sintering Finishing Alpha6 1.8 kilos 2.0 liters 0.20 hours 0.80 hours Zeta7 3.0 kilos 4.5 liters 0.35 hours 0.90 hours Information relating to materials purchased and materials used in production during May follows: Material Purchases Purchase Cost Standard Price Used in Production X442 14,500 kilos $52,200 $3.50 per kilo 8,500 kilos Y661 15,500 liters $20,925 $1.40 per liter 13,000 liters The following additional information is available: a. The company recognizes price variances when materials are purchased. b. The standard labor rate is $19.80 per hour in Sintering and $19.20 per hour in Finishing. c. During May, 1,200 direct labor-hours were worked in Sintering at a total labor cost of $27,000, and 2,850 direct labor-hours were worked in Finishing at a total labor cost of $59,850. d. Production during May was 1,500 Alpha6s and 2,000 Zeta7s. Requied: 1. Complete the standard cost card for each product, showing the standard cost of direct materials and direct labor. (Round your answers to 2 decimal places.) 2. Compute the materials price and quantity variances for each material. (Input all amounts as positive values. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e, zero variance).) 3. Compute the labor rate and efficiency variances for each operation. (Input all amounts as positive values. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e, zero variance).
Answer #1
STANDARD QUANTITY PER UNIT
PRODUCT
Alpha6 Zeta7
Raw materials:
X442(Kilos) 1.8 3.0
Y661(litres) 2.0 4.5
Sinterings (Labor hour) 0.20 0.35
Finishing (Labor hour) 0.80 0.90
DIRECT MATERIALS:
X442 Y661
A Materials Purchased 14500 15500
B Purchase Cost(Actual) $52,200 $20,925
C Standard Price $3.50 $1.40
D Quantity used in production (Actual) 8500 13000
E=B/A Actual Price paid $3.60 $1.35
DIRECT LABOR:
Sintering Finishing
A Standard Labor Rate(Per hour) $19.80 $19.20
B Direct labor hours (Actual) 1200 2850
C Total Labor Cost(Actual) $27,000 $59,850
D=C/B Actual Labor Rate per hour $22.50 $21.00
PRODUCT
Alpha6 Zeta7
Actual Quantity Produced 1500 2000
1 STANDARD COST CARD
ALPHA 6
Direct Material cost per unit $9.10 (X442:(1.8*3.5)+Y661:(1.4*2.0)
Direct Labor Cost Per Unit $19.32 (Sintering :(19.8*0.2)+Finishing :(19.2*0.8)
ZETA 7
Direct Material cost per unit $16.80 (X442:(3*3.5)+Y661:(1.4*4.5)
Direct Labor Cost Per Unit $24.21 (Sintering :(19.8*0.35)+Finishing :(19.2*0.9)
2 MATERIAL PRICE AND QUANTITY VARIANCE
SP1 Standard Price of X442 (per Kilo) $3.50
SP2 Standard Price of Y661 (per Liter) $1.40
AQ1 Actual Quantity of X442 Purchased 14500
AQ2 Actual Quantity of Y661 Purchased 15500
AP1 Actual Price of X442 (per Kilo) $3.60
AP2 Actual Price of Y661 (per Liter) $1.35
Material Price Variance X442=AQ1*(AP1-SP1) $1,450.00 (14500*(3.6-3.5)
Material Price Variance X442= $1,450.00 Unfavorable (Actual Price is higher than standard
Material Price Variance Y661=AQ2*(AP2-SP2) ($775.00)
Material Price Variance X442= $775.00 Favorable (Actual Price is Less than standard
Material Price Variance =AQ1*(AP1-SP1)+AQ2*(AP2-SP2)
Material Price Variance = $675.00 (14500*(3.6-3.5)+15500*(1.35-1.4)
Material Price Variance = $675.00 UNFAVORABLE
AQ3 Actual Quantity of X442 Used 8500
AQ4 Actual Quantity of Y661 Used 13000
SQ3 Standard Quantity of X442 for actual Production 8700 (1500*1.8+2000*3)
SQ4 Standard Quantity of Y661 for actual Production 12000 (1500*2.0+2000*4.5)
Material Quantity Variance X442=SP1*(AQ3-SQ3) $700.00 (3.5*(8500-8700)
Material Quantity Variance X442 $700.00 Favorable (Actual Quantity is Less than standard
Material Quantity Variance Y661=SP2*(AQ4-SQ4) $1,400.00 Unfavorable (Actual Quantity is more than standard
Material Quantity Variance :SP1*(AQ3-SQ3)+SP2*(AQ4-SQ4)
Material Quantity Variance = $700.00 (3.5*(8500-8700)+1.4*(13000-12000)
Material Quantity Variance = $700.00 Unfavorable
3 LABOR RATE AND EFFICIENCY VARIANCE Standard Labor Rate for Sintering $19.80 SR1 $19.20 Standard Labor Rate for Finishing SRLabor Rate Variance -SH1*(AR1-SR1)+SH2*(AR2-SR2) $8,100.00 (1000*(22.50-19.80)+3000*(21-19.20) $8,100.00 Labor Rate Variance